Find the equation of a line described as follows and express your answer in point-slope form, slope-intercept form, and standard form
Find the equation of the line parallel to the line y= 4/3x-6 that passes through the point (-6,2)

Answer by ewatrrr(24785)

point-slope form, y -2 = (4/3)(x+6)
slope-intercept form, y = (4/3)x + 10
and standard form 4x - 3y = -30